You are the extremely brilliant business statistician hired to analyze the data at Elliot's World o Tacos. Elliot wants to make sure that the high school kids he hired are putting the right amount of cheese (2oz) on the tacos. Too much cheese and he will lose money! Too little cheese and people complain. The tacos that should have 2 oz of cheese actually vary in amount with a standard deviation of 0.15 ounces. Assume the cheesiness of the tacos is normally distributed a. What percentage of all tacos made contains more than 2.22 ounces of cheese? b. What percentage of all tacos made contains less than 1.50 ounces? c. What percentage of tacos made contains between 1.46 to 2.38 ounces? d. Ninety-five percent of the tacos will contain at least how many ounces of cheesy goodness? e. What percentage of the tacos contains between 1.41 and 1.78 ounces?

PLEASE SHOW DETAILED WORK ON EACH ANSWER
Answer #1

Sar = 3.1 =O15 ( x 22.33) = 1- ( 3 4 22-3) 015 2 I-PL 2 47) 009292 OO7O8 외장 01/5 DP <150)=; < 50-3 =p(2.<-3.33) 0 16004 DPCI-